GriddPP PMB Meeting 825


Present: David Britton (Chair), Tony Cass, Peter Clarke, David Colling, Davide Costanzo, Tony Doyle  , Katy Ellis, Peter Gronbech, Jonathan Hays, Roger Jones, Andrew McNab Andrew Sansum, Sam Skipsey (Minutes), Jill Sambrook (Minutes)

Apologies: Alastair Dewhurst, David Kelsey, Steve Lloyd


  1. 1e6 Run [DB/All]

A discussion was held on an opportunity for collaboration with GridPP.

2) Quarterly report [SS]

Sam updated the group on the results and summary from quarterly report 09_Q12022.

Tier 1 report summary:

  • ALICE METRICS – lower usage was seen in quarter 1, but this was discussed at the experiment Liaison meeting and it was confirmed numbers and usage has since increased as we entered Run 3. The numbers should improve significantly for the next report.
  • WLCG CMS Service availability was at 86% instead of 98% target. WebDAV now working much better so this has now improved and should also show improvements next quarter. It was perhaps a small blip that has now been resolved at can report this at next OC meeting/report.
  • CMS efficiency @ Tier-1 has moved to amber, but is also similar to other Tier-1s, and should now (Q2) be up to 70%.  
  • DUNE also showing lower efficiency than metric.

AM said this should all improve when CPU scheduling is improved (by DUNE) and is not a worry at this time. This is something that is being worked through at the moment and should be resolved.

Tier 2’s:

  • A few areas around Argo availability discussed

Requests for Q2 reports has been sent out.

3) GridPP48 Ambleside [SS & JS]

Practical details:

  • JS has made contact with the venue and organisers. Arrangements around accommodation, parking, catering and final timetable to be confirmed by 11th of August. JS to continue to liaise with venue and organise this.
  • There are currently 54 registered attendees.
  • RJ to contact The Priest Hole to see if numbers can be increased for the conference dinner. (Will try and increase to 60).

Agenda and talks update:

  • Agenda is now nearly complete and SS to make this public soon. (Currently in draft mode).
  • SKA talk (Ian C)
  • IRIS talk – JH agreed he would be happy to give a 10 minutes session on IRIS if not enough elsewhere on the agenda. SS to get back to him this week.
  • LZ talk (needs speaker)
  • DUNE talk (AM)
  • Rubin/LSST talk. PC suggested Peter Love and/or Matt Doidge.

Sponsorship update:

  • No update from Alistair just yet, SS to email.
  • DB drafted a document to put on the website about sponsorship. It was agreed this can now be added by SS and link then sent to Andrew Sansum and Alistair.


SI-0 Tier-1 Manager’s weekly report & Technical Meetings [AD]
Not in attendance. No update this week.  

SI-1 ATLAS Weekly Review and Plans [DCos]
There have been a few smaller, sporadic issues over the past week.
Glasgow have been having issues with stage-in time outs.
Lancaster had a few failures over the weekend with VPN transfers, which are not fully understood.
These have now been resolved and nothing major to report.

SI-2 CMS Weekly Review and Plans [DC/KE]
A pretty smooth week over all.

  • gsiftp failing intermittently
    • LHCb submitted a ticket on this. One of the gateways had wrong keyring.
  • Tier 1 – batch farm scaled back on Tuesday due to heat as a precaution. Put back at 8pm once the immediate threat had passed and things started to cool down.
  • Looking into MINOS tapes. DB asked if we can get a summary of this once sorted and we can put a plan in place to learn from this moving forward. Look at this again at an upcoming meeting.

SI-3 LHCb Weekly Review and Plans [AM]

  • There was an issue with submission to the RAL PPD site on Friday morning, but issue seemed to be resolved later that day.
  • There still remains 3 open tickets. There have been some clarification comments, but no real progress. DB to check through tickets and then discuss at the Liaison meeting on Wednesday, however, there might not be much movement this week as Alistair is currently away.

SS can also ask at Xrootd dev meeting.

SI-4 Operations Meeting Report [SS,PG,PC] –
NTR that has not already been discussed.

SI-5 LCG Management Board Report of Issues [DB]
WLCG GMB meeting will take place next week (previous one delayed)

SI-6 External Contexts (eg NGI/EGI) [PC/JH]
JH – IRIS has received a preliminary report on funding. Should hear today if they get this money. If they do get it should be using some for digital assets.


800.5 AD – Arrange in person DIRAC/Rucio meeting at IC (Jan22). -ongoing

– Plans to re-establish virtual meetings and then set up an in person meeting. AD to ask Tim Noble if he can arrange this.

818.5 – DB/SS Group to formalise VOs to be added to the approved list. Ongoing

DB and SS have begun to look into this. 3 in particular were flagged as perhaps needing to be removed.
MaGrid (Morocco institute)
Raised with TBSupport for site feedback on who supports these and SS will wait until he has some feedback about use.
Discussion also around Microboone. 
Close to closing this off.

822.1 – AD to provide update about sponsorship for GridPP48 Ambleside.
SS to add web content about sponsorship to website and send to AD and AS.

823.1 – AD to Invite Rob to the Wednesday Liaison meeting for a few weeks to discuss LHCb tickets and leave as a standing item at the start of the agenda for a few weeks.

824.1 – SS will take lead on OsC action:
Collaboration to provide brief, narrative information on data recovery processes in the report for the next OsC meeting. – Ongoing

824.2 – DK and AD to draft a short narrative addressing OsC action:
Collaboration to provide narrative in the OsC report on the process that is used to determine the networking costs for GridPP6 for the next OsC meeting
DB asked for there to be a brief statement for GridPP6 with more details and plan for GridPP7.
DK and AD can send draft to DB. – Ongoing

824.3 – DB will take lead on OsC action:
Collaboration to request a user report from the ALICE experiment for all future OsC reports.
Some discussion around who was best to approach to provide this report. DB to contact someone at Birmingham and start the conversation about this. DB can then liaise with AD moving forward.

DB talked to Paul Newman at Brimingham and Mark. They are happy to do this. SS and DB to prod them at appropriate point ahead of next OsC.
Action now complete.

824.4 – JS and SS to now contact organisers in Ambleside to make practical arrangements, including trying to sort accommodation queries. – Complete

824.5 – DB and SS to create content calling for sponsors to be added to the website. Content will be circulated to the group for feedback before this is added to the website.
Content will be added to website today. Send link to AS and Alistair.  – Complete.

KE was asking about a significant activity on the GridPP VO @ RAL. Group discussed this and it appears it may be Giuseppe at Edinburgh for EUCLID. KE to liaise with PC.